a Square-based pyramid has a base of side 3.2m and a vertical height of 9.2m.
Find the length of the sloped edges of the pyramid

Answer by addingup(3677)

The base is 3.2. Draw a line from one corner to its opposite corner. Now you have two right triangles. Find the length of your line:
a^2+b^2 = c^2
3.2^2+3.2^2 = 20.48, take the square root: 4.525
The distance from the center of the triangle to this corner is 1/2 of that:
4.525/2 = 2.263
Now you have another right triangle, this one formed by the height 9.2, the length of the short leg which I just calculated, 2.263, and the sloped edge. In this triangle the sloped edge is the hypotenuse. Go find it:
2.263^2+9.2^2 = c^2 use your calculator and don't forget to take the square root of the sum of the squares to find the length of your corner. You should get 9.474, this is the length of the sloped edges.
:
John